# Positivity of Schur forms for Griffiths positive vector bundles of rank three over complex threefolds

### v1

@@ Abstract.
In this paper, we prove the positivity of the double mixed discriminant associated with a positive linear map between spaces of $3\times 3$ complex matrices, thereby settling the three-dimensional case of Finski’s open problem. As an application, we show that all Schur forms are weakly positive for Griffiths positive Hermitian holomorphic vector bundles of rank three over complex threefolds. This yields a complete affirmative answer, in the case where both the rank and the dimension are three, to the question posed by Griffiths in 1969.
@@
Positivity, Schur forms, low-rank, Griffiths positive vector bundle
@@ Introduction
Let $(E,h^{E})$ be a Hermitian holomorphic vector bundle of rank $r$ over a complex manifold $X$ of complex dimension $n$. Its Chern forms $c_{i}(E,h^{E})\in A^{i,i}(X)$ ($0\leq i\leq r$) are defined by the Chern–Weil formula
@@
where $R^{E}\in A^{1,1}(X,\mathrm{End}(E))$ denotes the Chern curvature of $(E,h^{E})$, and $c_{i}(E,h^{E})$ is the $(i,i)$-component of the total Chern form $c(E,h^{E})$. Fix an integer $k$ with $1\leq k\leq n$, and let $\lambda=(\lambda_{1},\dots,\lambda_{r})$ be a partition of $k$ (allowing trailing zeros) such that
@@
which is a closed real $(k,k)$-form (with the conventions $c_{0}=1$ and $c_{m}=0$ for $m<0$ or $m>r$). Schur forms recover the Chern forms and the signed Segre forms as special cases, and thus provide a unified family of characteristic forms naturally attached to $(E,h^{E})$. In particular, when $r=3$ and $k=3$, there are exactly three Schur forms, namely
@@
A fundamental theme in complex geometry is to understand how curvature positivity of a Hermitian holomorphic vector bundle controls positivity properties of its characteristic forms. On the numerical side, Bloch–Gieseker [BG71] proved numerical positivity of Chern classes for ample vector bundles, and Fulton–Lazarsfeld [FL83, Theorem I] showed that all Schur polynomials are numerically positive for ample vector bundles. Griffiths suggested that an analogous statement should hold at the level of differential forms for bundles with Griffiths positive curvature. More precisely, he conjectured that Schur forms are weakly positive in the sense of complex differential forms; see [FIN22, p. 1541, Question of Griffiths] for this formulation.
@@
Recall that a real $(k,k)$-form $u$ is weakly positive (resp. weakly nonnegative) if
@@
The conjecture has been verified in several important cases. Griffiths [GRI69, p. 249] proved positivity of the second Chern form for Griffiths positive bundles, and Güler [GUL12, Theorem 1.1] established Griffiths’ question for all signed Segre forms. Diverio–Fagioli [DF22] obtained weak positivity for a number of further polynomials in Chern forms via pushforward formulas on flag bundles; see also [FAG22a, FAG22b]. For stronger curvature positivity notions, much more is known. Bott–Chern [BC65, Lemma 5.3,(5.5)] proved nonnegativity of all Chern forms under Bott–Chern nonnegativity, and Li [LI21, Proposition 3.1] extended this to all Schur forms. Finski showed that Bott–Chern nonnegativity is equivalent to dual Nakano nonnegativity [FIN22, Theorem 2.15], proved nonnegativity of all Schur forms for Nakano nonnegative bundles by an algebraic argument [FIN22, Section 3.4], and obtained positivity for (dual) Nakano positive bundles via a differential-form refinement of the Kempf–Laksov formula [FIN22, Theorem 1.1]. More recently, the author [WAN24] proved positivity of Schur forms under the (weaker) assumption of strong decomposable positivity, again by algebraic methods.
A major insight of Finski [FIN22, Theorem 1.3] is that Griffiths’ Question 0.1 is equivalent to an operator-theoretic positivity problem involving mixed discriminants. Let $V$ be a complex vector space of dimension $r$. The mixed discriminant $\mathrm{D}_{V}:\mathrm{End}(V)^{\otimes r}\to\mathbb{C}$ is the polarization of the determinant introduced by Alexandrov [ALE39, Section 1]. In coordinates, for $A^{i}=(a^{i}_{k\bar{\ell}})_{k,\ell=1}^{r}$ ($i=1,\dots,r$), one has
$\mathrm{D}_{V}(A^{1},\dots,A^{r})=\frac{1}{r!}\frac{\partial^{r}}{\partial t^{1}\cdots\partial t^{r}}\det\bigl(t^{1}A^{1}+\cdots+t^{r}A^{r}\bigr)=\frac{1}{r!}\sum_{\sigma\in S_{r}}\det\bigl(a^{\sigma(i)}_{i\bar{k}}\bigr)_{i,k=1}^{r}.$
Using the natural identification $\mathrm{End}(V)\simeq\mathrm{End}(V)^{*}$, let $\mathrm{D}_{V}^{*}:\mathbb{C}\to\mathrm{End}(V)^{\otimes r}$ be the dual map. Finski posed the following question and proved that it is equivalent to Griffiths’ Question 0.1; see [FIN22, Theorem 1.3]. (For a fixed value of $r=\mathrm{rank}(E)=\dim V$, however, Problem 0.2 and Question 0.1 need not be equivalent.)
@@
In rank $3$, Fagioli [FAG22a, Theorem 0.1] proved weak nonnegativity of the Schur form
@@
for Griffiths semipositive bundles, and established the chain of inequalities
@@
Moreover, by twisting with a line bundle of sufficiently small negative curvature, one can upgrade this to weak positivity of $P_{(2,1,0)}\bigl(c(E,h^{E})\bigr)$ for Griffiths positive bundles of rank three (see Proposition 2.13). In addition, Güler [GUL12, Theorem 1.1] proved that all signed Segre forms are weakly positive. In particular, for rank-three bundles this yields weak positivity of the Schur forms corresponding to the partitions $(1,0,0)$, $(1,1,0)$, and $(1,1,1)$. Finally, the second Chern form $c_{2}(E,h^{E})=P_{(2,0,0)}\bigl(c(E,h^{E})\bigr)$ is also weakly positive; see [GRI69, FAG22a] and also Proposition 2.13.
Consequently, when both the rank and the complex dimension are three, Griffiths’ Question 0.1 reduces to proving weak positivity of the top Chern form $c_{3}(E,h^{E})=P_{(3,0,0)}(c(E,h^{E}))$. Combining this reduction with [FIN22, Theorem 3.4 and Corollary 3.5], we conclude that, in the case $r=\mathrm{rank}(E)=\dim V=3$ and $\dim X=n=3$, Problem 0.2 and Question 0.1 are equivalent.
In this paper we focus on Question 0.1 and Problem 0.2 in rank $3$ over a complex threefold (and, for comparison, we also discuss rank $2$). Write $H(E_{i\bar{j}})=B_{i\bar{j}}\in\mathrm{End}(W)$ for the images of the matrix units. Then the double mixed discriminant
@@
Following Gurvits [GUR04, Definition 4.2], one introduces the operator scaling
@@
where $C_{1}\in\mathrm{End}(W)$ and $C_{2}\in\mathrm{End}(V)$ are invertible. This scaling preserves positivity of the map and does not change the sign of $\Phi$. Moreover, by [GUR04, Theorem 4.7], one can choose invertible matrices $C_{1}$ and $C_{2}$ so that $S_{C_{1},C_{2}}(\frac{1}{r}H)$ becomes doubly stochastic. Consequently, without loss of generality we may work under the normalization
@@
Under this normalization, Problem 0.2 reduces to proving $\Phi>0$ for every family $\{B_{i\bar{j}}\}$ satisfying the Griffiths-type positivity condition (0.2).
Our main result gives an affirmative answer, in both rank and dimension are three, to both Griffiths’ Question 0.1 and Finski’s Problem 0.2.
@@
In particular, $\mathrm{D}_{W}\circ H^{\otimes\dim W}\circ\mathrm{D}_{V}^{*}\in\mathbb{R}$ is positive if $H:\mathrm{End}(V)\to\mathrm{End}(W)$ is a positive linear map and $\dim V=\dim W=3$.
Moreover, all Schur forms are weakly positive for Griffiths positive Hermitian holomorphic vector bundles of rank three over complex threefolds.
The article is organized as follows. In Section 1 we recall the notions of weak positivity for differential forms and Griffiths positivity for Hermitian holomorphic vector bundles. In Section 2 we reformulate Finski’s open problem as Open Problem 2.4 and solve it for $r=3$. In particular, Theorem 0.3 is proved in this section.
@@ 2.1. Finski’s open problem
$A_{p\bar{q}}:=(\theta_{i\bar{j},p\bar{q}})_{i,\bar{j}}\in\mathrm{End}(V),\qquad B_{i\bar{j}}:=(\theta_{i\bar{j},p\bar{q}})_{p,\bar{q}}\in\mathrm{End}(W).$
@@ 2.3. The case of rank three
$\mathrm{D}_{W}\circ H^{\otimes\dim W}\circ\mathrm{D}_{V}^{*}$

### v2

# Positivity of the third Chern form for Griffiths positive vector bundles

@@ Abstract.
In this paper, we prove the positivity of the double mixed discriminant associated with a positive linear map between spaces of third-order complex matrices, thereby settling the three-dimensional case of Finski’s open problem. As an application, we obtain the weak positivity of the third Chern form for Griffiths positive vector bundles. Moreover, we show that all Schur forms are weakly positive for Griffiths positive vector bundles of rank three over complex threefolds. This yields a complete affirmative answer, in the case where both the rank and the dimension are three, to the question posed by Griffiths in 1969.
@@
Positivity, Schur forms, the third Chern form, Griffiths positive vector bundle
@@ Introduction
A central problem in complex differential geometry is to understand how curvature positivity of a Hermitian holomorphic vector bundle constrains the positivity of its characteristic forms. Let $(E,h^{E})$ be a Hermitian holomorphic vector bundle of rank $r$ over a complex manifold $X$ of complex dimension $n$. Its Chern forms $c_{i}(E,h^{E})\in A^{i,i}(X)$ ($0\leq i\leq r$) are defined by
@@
where $R^{E}\in A^{1,1}(X,\mathrm{End}(E))$ denotes the Chern curvature of $(E,h^{E})$, and $c_{i}(E,h^{E})$ is the $(i,i)$-component of $c(E,h^{E})$.
Fix $1\leq k\leq n$, and let $\lambda=(\lambda_{1},\dots,\lambda_{r})$ be a partition of $k$ (allowing trailing zeros) with
@@
a closed real $(k,k)$-form (with the conventions $c_{0}=1$ and $c_{m}=0$ for $m<0$ or $m>r$). Schur forms simultaneously generalize Chern forms and (signed) Segre forms, and thus provide a unified family of characteristic forms naturally attached to $(E,h^{E})$. For instance, when $r=3$ and $k=3$, there are exactly three Schur forms:
@@
On the numerical side, Bloch–Gieseker [BG71] proved numerical positivity of Chern classes for ample vector bundles, and Fulton–Lazarsfeld [FL83, Theorem I] showed that all Schur polynomials are numerically positive for ample vector bundles. Griffiths proposed an analytic counterpart: for bundles with Griffiths positive curvature, the corresponding Schur forms should be (weakly) positive as differential forms. A convenient formulation is the following; see [FIN22, p. 1541, Question of Griffiths].
@@
We recall that a real $(k,k)$-form $u$ is weakly positive (resp. weakly nonnegative) if
@@
Griffiths [GRI69, p. 249] proved positivity of the second Chern form for Griffiths positive bundles, and Güler [GUL12, Theorem 1.1] established Griffiths’ question for all signed Segre forms. Diverio–Fagioli [DF22] obtained weak positivity for further polynomials in Chern forms via pushforward formulas on flag bundles; see also [FAG22a, FAG22b]. For stronger curvature positivity notions, more is known: Bott–Chern [BC65, Lemma 5.3,(5.5)] proved nonnegativity of all Chern forms under Bott–Chern nonnegativity, and Li [LI21, Proposition 3.1] extended this to all Schur forms. Finski showed that Bott–Chern nonnegativity is equivalent to dual Nakano nonnegativity [FIN22, Theorem 2.15], proved nonnegativity of all Schur forms for Nakano nonnegative bundles by an algebraic argument [FIN22, Section 3.4], and obtained positivity for (dual) Nakano positive bundles via a differential-form refinement of the Kempf–Laksov formula [FIN22, Theorem 1.1]. More recently, the author [WAN24] proved positivity of Schur forms under the (weaker) assumption of strong decomposable positivity.
A major insight of Finski [FIN22, Theorem 1.3] is that Griffiths’ Question 0.1 is equivalent to a structural positivity statement for positive linear maps expressed in terms of mixed discriminants. Let $V$ be a complex vector space of dimension $r$. The mixed discriminant $\mathrm{D}_{V}:\mathrm{End}(V)^{\otimes r}\to\mathbb{C}$ is the polarization of the determinant introduced by Alexandrov [ALE39, Section 1]. Using the natural identification $\mathrm{End}(V)\simeq\mathrm{End}(V)^{*}$, let $\mathrm{D}_{V}^{*}:\mathbb{C}\to\mathrm{End}(V)^{\otimes r}$ be the dual map. Finski posed the following question and proved that it is equivalent to Griffiths’ Question 0.1; see [FIN22, Theorem 1.3]. (For a fixed rank $r=\mathrm{rank}(E)=\dim V$, however, Problem 0.2 and Question 0.1 need not be equivalent.)
@@
In rank $3$, Fagioli [FAG22a, Theorem 0.1] proved weak nonnegativity of
@@
for Griffiths semipositive bundles, and established the inequalities
@@
By twisting with a line bundle of sufficiently small negative curvature, this can be upgraded to weak positivity of $P_{(2,1,0)}(c(E,h^{E}))$ for Griffiths positive bundles of rank three (see Proposition 2.13). In addition, Güler [GUL12, Theorem 1.1] proved that all signed Segre forms are weakly positive; in rank three this yields weak positivity of the Schur forms corresponding to $(1,0,0)$, $(1,1,0)$, and $(1,1,1)$. Finally, the second Chern form $c_{2}(E,h^{E})=P_{(2,0,0)}(c(E,h^{E}))$ is also weakly positive; see [GRI69, FAG22a] and Proposition 2.13. Consequently, when $\mathrm{rank}(E)=\dim X=3$, Griffiths’ Question 0.1 reduces to proving weak positivity of the top Chern form $c_{3}(E,h^{E})=P_{(3,0,0)}(c(E,h^{E}))$. Combining this reduction with [FIN22, Theorem 3.4 and Corollary 3.5], we conclude that, in the case $r=\mathrm{rank}(E)=\dim V=3$ and $\dim X=n=3$, Problem 0.2 and Question 0.1 are equivalent.
Main results. In this paper we settle the first nontrivial case $\dim V=\dim W=3$ of Finski’s Problem 0.2 and, equivalently, give a complete affirmative answer to Griffiths’ Question 0.1 for Griffiths positive bundles of rank $3$ on complex threefolds. Write $H(E_{i\bar{j}})=B_{i\bar{j}}\in\mathrm{End}(W)$ for the images of the matrix units. Then the double mixed discriminant
@@
Following Gurvits [GUR04, Definition 4.2], we consider the operator scaling
@@
where $C_{1}\in\mathrm{End}(W)$ and $C_{2}\in\mathrm{End}(V)$ are invertible. This scaling preserves positivity and does not change the sign of $\Phi$. Moreover, by [GUR04, Theorem 4.7], one can choose $C_{1},C_{2}$ so that $S_{C_{1},C_{2}}(\frac{1}{r}H)$ becomes doubly stochastic. Consequently, without loss of generality we work under the normalization
@@
Under (0.2) and (0.3), Problem 0.2 reduces to proving $\Phi>0$.
@@
In particular, $\mathrm{D}_{W}\circ H^{\otimes\dim W}\circ\mathrm{D}_{V}^{*}\in\mathbb{R}$ is positive if $H:\mathrm{End}(V)\to\mathrm{End}(W)$ is a positive linear map and $\dim V=\dim W=3$. Moreover, all Schur forms are weakly positive for Griffiths positive vector bundles of rank three over complex threefolds.
Remark 0.4.
By [FIN22, Proposition 2.21], the case $\dim V=\dim W=3$ is the first dimension in which a positive map is not necessarily decomposable in the sense of [FIN22, (2.40)] (as shown by Choi’s example [CHO75]). Thus Theorem 0.3 verifies Finski’s open problem in the first genuinely new dimension and provides a strong indication that the general statement may also hold.
In particular, Theorem 0.3 implies that the third Chern form $c_{3}(E,h^{E})$ is weakly positive when $(E,h^{E})\to X$ is Griffiths positive and $\mathrm{rank}(E)=\dim X=3$. We next extend this conclusion to arbitrary rank $r\geq 3$ on any complex manifold $X$ with $\dim X\geq 3$. Indeed, $c_{3}(E,h^{E})$ is the sum of all $3\times 3$ principal minors of the Chern curvature matrix $(R^{E})^{i}_{\,j}$:
(0.4) $c_{3}(E,h^{E})=\left(\tfrac{\sqrt{-1}}{2\pi}\right)^{3}\sum_{1\leq i_{1}<i_{2}<i_{3}\leq r}\det\bigl((R^{E})^{i}_{\,j}\bigr)_{i,j\in\{i_{1},i_{2},i_{3}\}}.$
Moreover, weak positivity of a real $(p,p)$-form can be tested by restricting to complex $p$-dimensional subspaces; see, for example, [FAG22a, Proposition 1.9]. Hence we may assume $\dim X=3$. Restricting $(E,h^{E})$ to the rank-three subbundle spanned (locally) by $\{e_{i_{1}},e_{i_{2}},e_{i_{3}}\}$ preserves Griffiths positivity, and applying Theorem 0.3 shows that each summand in (0.4) is weakly positive. Therefore their sum $c_{3}(E,h^{E})$ is weakly positive. We thus obtain:
Theorem 0.5.
Let $(E,h^{E})$ be a Griffiths positive vector bundle of rank $r\geq 3$ over a complex manifold $X$ with $\dim X\geq 3$. Then the third Chern form $c_{3}(E,h^{E})$ is weakly positive.
Remark 0.6.
Our result yields an elementary proof of the following fact: if $X$ is a compact complex threefold admitting a Hermitian metric with negative holomorphic bisectional curvature, then its Euler characteristic is negative. Indeed,
$\chi(X)=\int_{X}c_{3}(T_{X})=-\int_{X}c_{3}(T_{X}^{*}),$
and negative holomorphic bisectional curvature implies that the cotangent bundle $T_{X}^{*}$ is Griffiths positive. Hence $c_{3}(T_{X}^{*})$ is weakly positive, so $\int_{X}c_{3}(T_{X}^{*})>0$, and therefore $\chi(X)<0$.
Proof idea and organization of the paper
Proof idea.
The proof of Theorem 0.3 combines a normalization-by-scaling step with a geometric positivity argument for a mixed-discriminant expression. We first use operator scaling to reduce to the doubly stochastic normalization (0.3), which preserves positivity and the sign of $\Phi$. Under this normalization, $\Phi$ can be rewritten in terms of mixed discriminants of explicit third-order Hermitian matrices associated to $(B_{i\bar{j}})$. Introducing $C(\xi)=\sum_{i,j}B_{i\bar{j}}\,\xi^{i}\overline{\xi^{j}}$, a spherical moment identity expresses the relevant mixed-discriminant combinations as an average over $\xi\in S^{5}$ of a symmetric polynomial in $C(\xi)$. Diagonalizing $C(\xi)$ reduces the integrand to an inequality for symmetric polynomials of the eigenvalues of $C(\xi)$, which follows from a sharp Schur inequality. This yields $\Phi>0$. Finally, the equivalence of Finski’s problem with Griffiths’ question in the $(3,3)$-case translates $\Phi>0$ into weak positivity of all degree-$3$ Schur forms, and Theorem 0.5 follows by restriction to rank-$3$ subbundles and the principal-minor formula (0.4).
Organization of the paper.
In Section 1 we recall weak positivity of differential forms and Griffiths positivity for Hermitian holomorphic vector bundles. In Section 2 we reformulate Finski’s open problem as Open Problem 2.4 and solve it for $r=3$; in particular, we prove Theorem 0.3.
Acknowledgments. The author is grateful to Siarhei Finski for helpful discussions concerning Remark 0.4, and to Vamsi Pritham Pingali for helpful discussions related to Theorem 0.5 and Remark 0.6.
@@ 2.1. Finski’s open problem
$A_{p\bar{q}}:=(\theta_{i\bar{j},p\bar{q}})_{i\bar{j}}\in\mathrm{End}(V),\qquad B_{i\bar{j}}:=(\theta_{i\bar{j},p\bar{q}})_{p\bar{q}}\in\mathrm{End}(W).$
@@ 2.3. The case of rank three
$\mathrm{D}_{W}\circ H^{\otimes\dim W}\circ\mathrm{D}_{V}^{*}(1)$